OpenStudy (anonymous):

1. Simplify: √48x^3y^2 2. Simplify: √144x + √36x - √25x 3. Simplify: √18x^3 / √3xy ____ 4. Solve for x: √3x+1=5

OpenStudy (anonymous):

@TrashMan914

OpenStudy (anonymous):

2.Simplify: Sqrt(144x) + Sqrt(36x) - Sqrt(25x) √144x + √36x - √25x = (12 + 6 + 5)√x = 23√x. 3.Simplify: Sqrt(18x^3) / Sqrt(3xy) √(18x^3) / √(3xy) = √ (18x^3 / 3xy) = √(6x^2 / y) = x√(6/y). 4. Solve for x: Sqrt(3x+1)=5 3x + 1 = 25 3x = 24 x = 8.

OpenStudy (anonymous):

@KunoichiCassy1234 We are not to work out the entire problem, nor are we to simply give them the answer to the questoin. It's against CoC: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

sorry I apologize

OpenStudy (anonymous):

I get what you're saying. So can you help me know how to work out numbers 1 and 5? You dont have to give me the answer?? I just need help understanding

OpenStudy (anonymous):

Hence the title, OpenStudy.

OpenStudy (anonymous):

>.> okay.